# Sharp decay estimates and numerical analysis for weakly coupled systems of two subdiffusion equations

### v1

@@ 4. Numerical schemes and stability analysis
$&\partial_{t}^{\alpha}(u-u_{0})-d_{1}u_{xx}+c_{11}(x,t)u+c_{12}(x,t)v=F_{1}(x,t),\\ &\partial_{t}^{\beta}(v-v_{0})-d_{2}v_{xx}+c_{21}(x,t)u+c_{22}(x,t)v=F_{2}(x,t)\end{aligned}&\mbox{in }(0,L)\times(0,T),\\ u=v=0&\mbox{on }\{0,L\}\times(0,T),\end{cases}$ (4.1)
where $L,d_{1},d_{2}$ are positive constants. In other words, we restrict the spatial dimension $d=1$ and the principal coefficients $A(x),B(x)$ as constants. Meanwhile, we allow the appearance of inhomogeneous terms $F_{1}(x,t),F_{2}(x,t)$ for the completeness of numerical schemes, which will be set as zero in the study of the asymptotic behavior. Also, we only discuss the case of $\alpha<1$, i.e., both orders are fractional, as that of $\alpha=1$ is classical.
@@
For the numerical computation, we assume that all functions $c_{k\ell},F_{k}$ allow pointwise definition at grid points.
Now we are well prepared to discretize the problem (4.1). Assume that we have finished the simulation of $u,v$ until $t=t_{n}$ and are in a position to proceed to the next level $t=t_{n+1}$ for some $n=0,\dots,N-1$. Regarding the lower order parts $-c_{k1}u-c_{k2}v+F_{k}$ ($k=1,2$) as given data at the previous level $t=t_{n}$, we propose a semi-implicit scheme for (4.1) as
@@
Following the same line, we regard the lower order parts $-c_{k1}u-c_{k2}v+F_{k}$ ($k=1,2$) as unknowns at the next level $t=t_{n+1}$ to propose a fully implicit scheme for (4.1) as
@@
[…]
Then we can reformulate (4.4) as a linear system
@@
$\sigma(\mathrm{A})\subset\bigcup_{i=1}^{2(I-1)}\mathrm{R}_{i},$
@@
Similarly, for $\mathrm{R}_{1},\mathrm{R}_{I-1},\mathrm{R}_{I},\mathrm{R}_{2(I-1)}$, we have
@@
$12r_{1}+|\gamma^{1}_{i}|-|\alpha_{i}|\geq 1,\quad 2r_{2}+|\gamma^{2}_{i}|-|\beta_{i}|\geq 1,\quad i=1,\dots,I-$ 1
@@
$&\partial_{t}^{\alpha}(u-u_{0})-d_{1}u_{xx}=G_{1}(u,v),\\ &\partial_{t}^{\beta}(v-v_{0})-d_{2}v_{xx}=G_{2}(u,v)\end{aligned}&\mbox{in }(0,L)\times(0,T),\\ u=v=0&\mbox{on }\{0,L\}\times(0,T),\end{cases}$
@@
Finally, we mention that all above discussions automatically work for more components than $2$, e.g., in the next section we will also deal with coupled systems with $3$ components. Likewise, the coupling of subdiffusion and usual diffusion equations can also be discretized in the same manner, and the absolute stability still holds as long as we keep the scheme implicit. The arguments are almost identical and we omit the details here.
@@ 5. Numerical verification of decay rates
and the source terms as $F_{1}=F_{2}=0$. As for the initial values, we are concerned with whether $v_{0}$ vanishes identically or not. In numerical examples, we consider the following cases:
$\mbox{(i)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=\frac{\pi}{2}-\left|x-\frac{\pi}{2}\right|,\\ \mbox{(ii)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=0.\end{aligned}\right.$ (5.1)
@@
$\mbox{(i)}\ &\mbox{If $u_{0}\not\equiv 0$ and $v_{0}\not\equiv 0$, then }\|u(t)\|,\|v(t)\|\sim t^{-0.5},\\ \mbox{(ii)}\ &\mbox{If $u_{0}\not\equiv 0$ and $v_{0}\equiv 0$, then }\|u(t)\|,\|v(t)\|\sim t^{-0.9}\end{aligned}\right.\quad\mbox{as }t\to+\infty.$
@@
Next, we study the more interesting case of $\alpha=1$ in (4.1) and change several different values of $\beta<1$. We first fix $\beta=0.5$ and repeat the same test in Example 5.1. From the numerical results plotted in Figure 3, we clearly see
$\mbox{(i)}\ &\mbox{If $u_{0}\not\equiv 0$ and $v_{0}\not\equiv 0$, then }\|u(t)\|,\|v(t)\|\sim t^{-0.5},\\ \mbox{(ii)}\ &\mbox{If $u_{0}\not\equiv 0$ and $v_{0}\equiv 0$, then }\|u(t)\|,\|v(t)\|\sim t^{-1.5}\end{aligned}\right.\quad\mbox{as }t\to+\infty.$
@@
$&\mbox{If $\alpha=1.0$ and $\beta=0.3$, then }\|u(t)\|,\|v(t)\|\sim t^{-1.3},\\ &\mbox{If $\alpha=1.0$ and $\beta=0.7$, then }\|u(t)\|,\|v(t)\|\sim t^{-1.7}\end{aligned}\right.\quad\mbox{as }t\to+\infty.$
@@
$&\partial_{t}^{\alpha}(u-u_{0})-u_{xx}+u-0.5v-0.5w=0,\\ &\partial_{t}^{\beta}(v-v_{0})-v_{xx}-0.5u+v-0.5w=0,\\ &\partial_{t}^{\gamma}(w-w_{0})-w_{xx}-0.5u-0.5v+w=0\end{aligned}&\mbox{in }(0,\pi)\times(0,T),\\ u=v=w=0&\mbox{on }\{0,\pi\}\times(0,T),\end{cases}$ (5.3)
where $01\geq\alpha\geq\beta\geq\gamma>$ 0. The coupling coefficients $c_{k\ell}$ are safely chosen to achieve both the numerical stability and the possible assumption for the expected decay rate. For 3 components, the combinations of initial values are more flexible than before and here we consider the following 3 cases:
$\mbox{(i)}\ &u_{0}(x)=x(\pi-x),\ v_{0}(x)=\sin x,\ w_{0}(x)=\frac{\pi}{2}-\left|x-\frac{\pi}{2}\right|,\\ \mbox{(ii)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=\frac{\pi}{2}-\left|x-\frac{\pi}{2}\right|,\ w_{0}(x)=0,\\ \mbox{(iii)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=w_{0}(x)=0.\end{aligned}\right.$ (5.4)
@@
$\mbox{(i)}\ &\mbox{If $u_{0}\not\equiv 0$, $v_{0}\not\equiv 0$ and $w_{0}\not\equiv 0$, then }\|u(t)\|,\|v(t)\|,\|w(t)\|\sim t^{-0.3},\\ \mbox{(ii)}\ &\mbox{If $u_{0}\not\equiv 0$, $v_{0}\not\equiv 0$ and $w_{0}\equiv 0$, then }\|u(t)\|,\|v(t)\|,\|w(t)\|\sim t^{-0.5},\\ \mbox{(iii)}\ &\mbox{If $u_{0}\not\equiv 0$, $v_{0}\equiv 0$ and $w_{0}\equiv 0$, then }\|u(t)\|,\|v(t)\|,\|w(t)\|\sim t^{-0.9}\end{aligned}\right.\quad\mbox{as }t\to+\infty.$
@@
We consider Cases (ii) in (5.4), namely, $w_{0}\equiv 0$ and $v_{0}\not\equiv 0$. Since now we require $\alpha=1$ and $\gamma<1$, we have some flexibility of choosing $\beta$, i.e., $\beta<1$ or $\beta=1$. We implement both cases and demonstrate the decay of numerical solutions in Figure 6, from which we clearly see
$&\mbox{If $\alpha=1.0$, $\beta=0.5$ and $\gamma=0.3$, then }\|u(t)\|,\|v(t)\|,\|w(t)\|\sim t^{-0.5},\\ &\mbox{If $\alpha=1.0$, $\beta=1.0$ and $\gamma=0.5$, then }\|u(t)\|,\|v(t)\|,\|w(t)\|\sim t^{-1.5}\end{aligned}\right.\quad\mbox{as }t\to+\infty.$
@@
Finally, we consider Case (iii) in (5.4), namely, $v_{0}=w_{0}\equiv 0$ and $u_{0}\not\equiv 0$. As before, we test both cases of $\beta<1$ and $\beta=1$ and illustrate the decay of numerical solutions in Figure 7. Contrary to the last example, here both cases show a decay rate of $t^{-1.5}$ independent of the choice of $\beta$.
@@
Until now, we have collected sufficient hints from numerical experiments to propose a general conjecture on the sharp long-time decay rates for coupled subdiffusion systems with more than 2 components.
@@
$&\partial_{t}^{\alpha_{k}}(u_{k}-u_{0}^{(k)})-\mathrm{div}(\bm{A}_{k}(\bm{x})\nabla u_{k})+\sum_{\ell=1}^{K}c_{k\ell}(\bm{x},t)u_{\ell}=0&\quad&\mbox{in }\Omega\times\mathbb{R}_{+},\\ &u_{k}=0&\quad&\mbox{on }\partial\Omega\times\mathbb{R}_{+},\end{aligned}\right.\quad k=1,\dots,K,$ (5.5)
where $1\geq\alpha_{1}\geq\cdots\geq\alpha_{K}>0,$ $\alpha_{K}<1$ and the coefficients $\bm{A}_{k},c_{k\ell}$ $(k,\ell=1,\dots,K)$ fulfill similar conditions as (1.2)–(1.3). Let $\underline{\alpha}$ be the lowest order among $\alpha_{1},\dots,\alpha_{K}$ whose corresponding initial value satisfies $u_{0}^{(k)}\not\equiv 0$. Under some technical assumption such as ($,$ there exists a constant $C>0$ such that the solution to (5.5) satisfy
@@ Appendix A Explicit decay rate of a decoupled subdiffusion system
In the previous section, we described the motivation of discovering Theorem 1.1 by means of numerical simulation. As a supplementary clue, in this appendix we provide another theoretical approach to finding the superlinear decay rate $t^{-(1+\beta)}$ of the solution to a special decoupled system of a usual diffusion equation and a subdiffusion one.
@@
where $B(\,\cdot\,,\,\cdot\,)$ denotes the Beta function. Then we substitute $Q(t,j)$ back into $P(t)$ to obtain
@@
Now we concentrate on the series $R(t)$. Rearranging the terms in $R(t)$ according to the power of $-\lambda_{n}t^{\beta}$, we recall the definition of the Mittag-Leffler functions to calculate
@@
=$=\left\{\frac{1}{\Gamma(\beta+1)}+\frac{(-\lambda_{n}t)}{\Gamma(\beta+2)}+\frac{(-\lambda_{n}t)^{2}}{\Gamma(\beta+3)}+\frac{(-\lambda_{n}t)^{3}}{\Gamma(\beta+4)}+\cdots\right\}$
@@
Then we plug the above expression of $R(t)$ back into $P(t)$ and then into $v(t)$ to represent
@@
Now we invoke the following asymptotic behavior for the Mittag-Leffler function $E_{\eta,\mu}(-z)$ with $\eta\in(0,2)$, $\mu>0$ and $z<0$ (see Podlubny [16, Theorem 1.4]):
@@
$-\Gamma(-\beta)>0,\quad O(t^{\beta-3})=o(t^{-(1+\beta)}),\ t\gg 1$
@@
where $C>0$ is a constant depending only on $\mathcal{A}$. The proof is completed. ∎
Owing to the simplicity of (A.1), we can easily write down its explicit solution, so that the above lemma provides not only the sharp decay rate of $v$ but also its limit pattern $u_{\infty}$, i.e., the profile of $v$ approaches $u_{\infty}$ as $t\to+\infty$. In other words, $v$ asymptotically takes the form of separated variables for large $t$ with the spatial component $\frac{u_{\infty}}{-\Gamma(-\beta)}$ and the temporal component $t^{-(1+\beta)}$. The limit pattern in more general situations is complicated, which can be another interesting future topic.
Acknowledgements
The first author thanks the National Natural Science Foundation of China (12271277), and Ningbo Youth Leading Talent Project (2024QL045). The second author is supported by JSPS KAKENHI Grant Numbers JP22K13954, JP23KK0049 and Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515012248). This work is partially supported by the Open Research Fund of Key Laboratory of Nonlinear Analysis & Applications (Central China Normal University), Ministry of Education, China.

### v2

@@ 4. Numerical schemes and stability analysis
$\begin{cases}\begin{aligned} &\partial_{t}^{\alpha}(u-u_{0})-d_{1}u_{xx}+c_{11}(x,t)u+c_{12}(x,t)v=F_{1}(x,t),\\ &\partial_{t}^{\beta}(v-v_{0})-d_{2}v_{xx}+c_{21}(x,t)u+c_{22}(x,t)v=F_{2}(x,t)\end{aligned}&\mbox{in }(0,L)\times(0,T),\\ u=v=0&\mbox{on }\{0,L\}\times(0,T),\end{cases}$ (4.1)
where $L,d_{1},d_{2}$ are positive constants. In other words, we restrict the spatial dimension $d=1$ and the principal coefficients $A(x),B(x)$ as constants. Meanwhile, the appearance of inhomogeneous terms $F_{1}(x,t),F_{2}(x,t)$ is allowed for the completeness of the numerical schemes, which will be set as zero in the study of asymptotic behaviors. Also, the discussion is limited to the case of $\alpha<1$, that is, both orders are fractional, as that of $\alpha=1$ is classical.
@@
For the numerical computation, all functions $c_{k\ell},F_{k}$ are assumed to allow pointwise definitions at grid points.
Now we are well prepared to discretize the problem (4.1). Assume that we have finished the simulation of $u,v$ until $t=t_{n}$ and are in a position to proceed to the next level $t=t_{n+1}$ for some $n=0,\dots,N-1$. Regarding the lower order parts $c_{k1}u-c_{k2}v+F_{k}$ ($k=1,2$) as given data at the previous level $t=t_{n}$, we propose a semi-implicit scheme for (4.1) as
@@
Following the same line, we regard the lower order parts $c_{k1}u-c_{k2}v+F_{k}$ ($k=1,2$) as unknowns at the next level $t=t_{n+1}$ to propose a fully implicit scheme for (4.1) as
@@
$\begin{aligned} \alpha_{i}&:=1+2r_{1}+\frac{\Delta x^{2}r_{1}}{d_{1}}c_{11}^{i,n+1},&\quad\gamma^{1}_{i}&:=\frac{\Delta x^{2}r_{1}}{d_{1}}c_{12}^{i,n+1},\\ \beta_{i}&:=1+2r_{2}-\frac{\Delta x^{2}r_{2}}{d_{2}}c_{22}^{i,n+1},&\quad\gamma^{2}_{i}&:=\frac{\Delta x^{2}r_{2}}{d_{2}}c_{21}^{i,n+1}\end{aligned}\quad(i=1,\dots,I-1).$
Then (4.4) can be reformulated as a linear system
@@
$\sigma(\mathrm{A})\subset\bigcup_{i=1}^{2(I-1)}\mathrm{R}_{i},\$
@@
Similarly, for $\mathrm{R}_{1},\mathrm{R}_{I-1},\mathrm{R}_{I},\mathrm{R}_{2(I-1)}$, it can be shown that
@@
$2r_{1}+|\gamma^{1}_{i}|-|\alpha_{i}|\geq 1,\quad 2r_{2}+|\gamma^{2}_{i}|-|\beta_{i}|\geq 1,\quad i=1,\dots,I-1$
@@
$\begin{cases}\begin{aligned} &\partial_{t}^{\alpha}(u-u_{0})-d_{1}u_{xx}=G_{1}(u,v),\\ &\partial_{t}^{\beta}(v-v_{0})-d_{2}v_{xx}=G_{2}(u,v)\end{aligned}&\mbox{in }(0,L)\times(0,T),\\ u=v=0&\mbox{on }\{0,L\}\times(0,T),\end{cases}$
@@
Finally, we mention that all above discussions automatically work for more components than $2$, e.g., in the next section we will also deal with coupled systems with $3$ components. Likewise, the coupling of subdiffusion and usual diffusion equations can also be discretized in the same manner, and the absolute stability still holds as long as we keep the scheme implicit. The arguments are almost identical and the details here are omitted.
@@ 5. Numerical verification of decay rates
and the source terms as $F_{1}=F_{2}=0$. As for the initial values, we are concerned with whether $v_{0}$ vanishes identically or not. In numerical examples, the following cases are considered.
$\left\{\!\begin{aligned} \mbox{(i)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=\frac{\pi}{2}-\left|x-\frac{\pi}{2}\right|,\\ \mbox{(ii)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=0.\end{aligned}\right$. (5.1)
@@
[…]
@@
Next, we study the more interesting case of $\alpha=1$ in (4.1) and change several different values of $\beta<1$. We first fix $\beta=0.5$ and repeat the same test in Example 5.1. From the numerical results plotted in Figure 3, it can be clearly observed that
[…]
@@
[…]
@@
$\begin{cases}\begin{aligned} &\partial_{t}^{\alpha}(u-u_{0})-u_{xx}+u-0.5v-0.5w=0,\\ &\partial_{t}^{\beta}(v-v_{0})-v_{xx}-0.5u+v-0.5w=0,\\ &\partial_{t}^{\gamma}(w-w_{0})-w_{xx}-0.5u-0.5v+w=0\end{aligned}&\mbox{in }(0,\pi)\times(0,T),\\ u=v=w=0&\mbox{on }\{0,\pi\}\times(0,T),\end{cases}$ (5.3)
where $1\geq\alpha\geq\beta\geq\gamma>0$. The coupling coefficients $c_{k\ell}$ are safely chosen to achieve both the numerical stability and the possible assumption for the expected decay rate. For 3 components, the combinations of initial values are more flexible than before and here we consider the following 3 cases:
$\left\{\!\begin{aligned} \mbox{(i)}\ &u_{0}(x)=x(\pi-x),\ v_{0}(x)=\sin x,\ w_{0}(x)=\frac{\pi}{2}-\left|x-\frac{\pi}{2}\right|,\\ \mbox{(ii)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=\frac{\pi}{2}-\left|x-\frac{\pi}{2}\right|,\ w_{0}(x)=0,\\ \mbox{(iii)}\ &u_{0}(x)=\sin x,\ v_{0}(x)=w_{0}(x)=0.\end{aligned}\right$. (5.4)
@@
[…]
@@
Consider Cases (ii) in (5.4), namely, $w_{0}\equiv 0$ and $v_{0}\not\equiv 0$. Since now we require $\alpha=1$ and $\gamma<1$, we have some flexibility of choosing $\beta$, i.e., $\beta<1$ or $\beta=1$. We implement both cases and demonstrate the decay of numerical solutions in Figure 6, from which it follows that
[…]
@@
Finally, Case (iii) in (5.4) is considered, namely, $v_{0}=w_{0}\equiv 0$ and $u_{0}\not\equiv 0$. As before, we test both cases of $\beta<1$ and $\beta=1$ and illustrate the decay of numerical solutions in Figure 7. Contrary to the last example, here both cases show a decay rate of $t^{-1.5}$ independent of the choice of $\beta$.
@@
6. Concluding remarks
This research established a fundamental dichotomy in the long-time decay behavior of weakly coupled subdiffusion systems provided that the initial value of the lower-order component vanishes identically. When the highest fractional order satisfies $0<\alpha<1$, the solutions exhibit characteristic sublinear decay $t^{-\alpha}$, consistent with the established pattern for the single fractional diffusion equation. Remarkably, systems with the highest order $\alpha=1$ demonstrate a qualitatively distinct superlinear decay pattern $t^{-(1+\beta)}$. This accelerated decay phenomenon fundamentally distinguishes coupled systems from single equations, where such a superlinear decay pattern never occurs.
As for the open problems related to the asymptotics for the coupled subdiffusion system, let us mention the decay patterns for
  1. (i)
    multi-component fractional differential systems, i.e., $K\geq 3$, and
  2. (ii)
    coupled time-fractional wave systems, i.e., the highest order is in $(1,2)$.
For (i), in the previous section, we have collected sufficient hints from numerical experiments to propose a general conjecture on the sharp long-time decay rates for coupled subdiffusion systems with more than $2$ components.
@@ 6. Concluding remarks
$\left\{\begin{aligned} &\partial_{t}^{\alpha_{k}}(u_{k}-u_{0}^{(k)})-\mathrm{div}(\bm{A}_{k}(\bm{x})\nabla u_{k})+\sum_{\ell=1}^{K}c_{k\ell}(\bm{x},t)u_{\ell}=0&\quad&\mbox{in }\Omega\times\mathbb{R}_{+},\\ &u_{k}=0&\quad&\mbox{on }\partial\Omega\times\mathbb{R}_{+},\end{aligned}\right$. (6.1)
[…]
@@
For (ii), the case $1<\alpha<2$ is indeed an important and interesting direction. In our current work, the decay pattern was established by employing the coercivity and maximum principle for fractional differential equations to derive our main results. However, for the case $1<\alpha<2$, the existing theoretical tools mentioned above are not yet well-established, and there are even known counterexamples. For instance, it was demonstrated in [6] that solutions of fractional differential equations in this range may exhibit finite oscillation behavior and have finitely many zeros, which fundamentally differs from the case $0<\alpha<1$, see e.g., as shown in the paper [13].
We plan to further investigate the properties of solutions in these cases and explore potential extensions of mathematical tools in subsequent research.
@@ Appendix A Explicit decay rate of a decoupled subdiffusion system
In Section 5, the motivation of discovering Theorem 1.1 by means of numerical simulation was described. As a supplementary clue, in this appendix we provide another theoretical approach to finding the superlinear decay rate $t^{-(1+\beta)}$ of the solution to a special decoupled system of a usual diffusion equation and a subdiffusion one.
@@
where B(,)B($\cdot\,$$\cdot\,$) denotes the Beta function. Then substituting $Q(t,j)$ back into $P(t)$ implies
@@
Now we concentrate on the series $R(t)$. Rearranging the terms in $R(t)$ according to the power of $\lambda_{n}t^{\beta}$, we recall the definition of the Mittag-Leffler functions to calculate
@@
$=\left\{\frac{1}{\Gamma(\beta+1)}+\frac{(-\lambda_{n}t)}{\Gamma(\beta+2)}+\frac{(-\lambda_{n}t)^{2}}{\Gamma(\beta+3)}+\frac{(-\lambda_{n}t)^{3}}{\Gamma(\beta+4)}+\cdots\right\}$
@@
Then plugging the above expression of $R(t)$ back into $P(t)$ and then into $v(t)$ yields
@@
Now we invoke the following asymptotic behavior for the Mittag-Leffler function $E_{\eta,\mu}(-z)$ with $\eta\in(0,2)$, $\mu>0$ and $z<0$ (see Podlubny [17, Theorem 1.4]):
@@
$\Gamma(-\beta)>0,\quad O(t^{\beta-3})=o(t^{-(1+\beta)}),\ t\gg 1$
@@
where $C>0$ is a constant depending only on $\mathcal{A}$. The proof is complete. ∎
Owing to the simplicity of (A.1), we can easily write down its explicit solution, so that the above lemma provides not only the sharp decay rate of $v$ but also its limit pattern $u_{\infty}$, i.e., the profile of $v$ approaches $u_{\infty}$ as $t\to+\infty$. In other words, $v$ asymptotically takes the form of separated variables for large $t$ with the spatial component $\frac{u_{\infty}}{-\Gamma(-\beta)}$ and the temporal component $t^{-(1+\beta)}$.
Moreover, based on the decay estimates, we demonstrate that if the solution decays faster than $t^{-(1+\beta)}$, then its initial value $u_{0}$ must be zero.
Corollary A.1.
Let $(u,v)$ be the solution to the fractional diffusion system (A.1) with initial data $u(x,0)=u_{0}(x)$ and $v(x,0)=0$. If there exists a constant $C>0$ such that
$\|v(t)\|_{H^{6}(\Omega)}\leq Ct^{-(1+\beta)}\eta(t)$
with an infinitesimal $\eta(t)$ as $t\to\infty$. Then the initial value $u_{0}$ must be zero.
The proof is trivial and is omitted. This corollary implies that any non-trivial initial data cannot yield a solution that decays at a rate exceeding the critical rate in Lemma A.1. A similar result for the single fractional diffusion-wave equation can be found in e.g., Yamamoto [22].
The limit pattern in more general situations is complicated, which can be another interesting future topic.
Acknowledgments
The authors thank the anonymous referees for valuable comments. The first author thanks the National Natural Science Foundation of China (12271277), and Ningbo Youth Leading Talent Project (2024QL045). The second author is supported by JSPS KAKENHI Grant Numbers JP22K13954, JP23KK0049 and Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515012248).